OpenStudy (anonymous):

6x+9/15x^2 divided by 16x-12/10x^4 options: a) x^2(2x+3)/2(4x-3) b)2x+3/2(4x-3) c)2(4x-3)/x^2(2x+3) d) 8x^2(4x-3)/9(2x+3)

OpenStudy (anonymous):

multiply by the reciprocal, which is the same as division \[\frac{6x+9}{15x^2}*\frac{10x^4}{16x-12}\] \[\frac{3(2x+3)*2x^2}{3*4(4x-3)}=\frac{6x^2(2x+3)}{12(4x-3)}=\frac{x^2(2x+3)}{2(4x-3)}\]

OpenStudy (anonymous):

thank you so so much !